The length of Mrs. Puff's motorboat is 60 ft. Mr. Krabs made a scale model of the same boat a few days later. The length of the scale model is just 9 in. What is the scale factor used to downsize Mrs. Puff's motorboat to create Mr. Krabs scale-model boat?
Answer by ikleyn(52908)

.
The length of Mrs. Puff's motorboat is 60 ft. Mr. Krabs made a scale model of the same boat a few days later.
The length of the scale model is just 9 in. What is the scale factor used to downsize Mrs. Puff's
motorboat to create Mr. Krabs scale-model boat?
~~~~~~~~~~~~~~~~~~~~~~~~~

Convert 60 feet to inches by multiplying by 12.


So,  60 ft = 60*12 inches = 720 inches.


The ratio 720/9 = 80.


So, the model is 1/80 of the real size.